OpenStudy (anonymous):

80f'(x)+69.67218*((e^(-0.00012005x))/((42/16.5)x+147))*f^2(x)=784

OpenStudy (anonymous):

Please help

OpenStudy (unklerhaukus):

\[80f'(x)+69.67218\times\frac{e^{-0.00012005x}}{\frac{42}{16.5}x+147}\times f^{2}(x)=784\] ?

OpenStudy (anonymous):

Yes

OpenStudy (unklerhaukus):

solve for x?

OpenStudy (anonymous):

For f(x)